# Optimal Control of Nonlinear Systems with Unknown Dynamics

### v2

@@ 6.4 Proof of Theorem 1
$+\hat{J}_{\mu}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*}).$
By selecting the step sizes as $\alpha_{k}^{f}=\frac{\beta_{f}}{2+k}$ from Lemma 1, $\alpha_{J}=(1-\gamma)/4$ from Lemma 3, we derive the following results:
$\sum_{k=0}^{K-1}\alpha_{k}^{\mu}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}-\frac{\sum_{k=0}^{K-1}(\alpha_{k}^{\mu})^{2}L_{J\theta}^{2}L_{J\theta\theta}}{2}$ (37)
$\leq$ $\hat{J}_{\mu}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})+C_{1}\Big{(}(1-\nu_{J})^{K/2}$
×$\times\parallel\boldsymbol{\theta}_{0}^{J}-\boldsymbol{\theta}^{J*}\parallel+\sqrt{\frac{\nu_{f}}{2+K}}+2\sum_{k=0}^{K-1}(\sqrt{\frac{\nu_{f}}{2+k+1}}$
@@
Furthermore, by applying the geometric series formula ($\sum_{k=0}^{\infty}ar^{k}=\frac{a}{1-r},|r|<1$) to (37), and subsequently taking the minimum of the left-hand side of (37), we obtain:
$\min_{k\in\{0,1,2,\cdots,K-1\}}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}$ (38)
@@
$+\sqrt{\nu_{f}}(K^{-1/2}+4K^{1/2})+\frac{1}{C_{1}}(\hat{J}_{\mu}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})$
$-\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*}))\Big{)}+\frac{\sum_{k=0}^{K-1}(\alpha_{k}^{\mu})^{2}L_{J\theta}^{2}L_{J\theta\theta}}{2\sum_{k=0}^{K-1}\alpha_{k}^{\mu}}.$
[…]
6.5 Proof of Corollary 2
@@
$\min_{k\in\{0,1,2,\cdots,K-1\}}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}$
@@
$+\hat{J}_{\mu}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})\Big{)}$
@@
$\quad+\hat{J}_{\mu}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})\Big{)},$

### v3

@@ 6.2.2 Convergence Analysis of the Estimated Policy Gradient
$+\hat{J}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*}).$
By choosing the step sizes as $\alpha_{k}^{f}=\frac{\beta_{f}}{2+k}$ (Lemma 1) and $\alpha_{J}=(1-\gamma)/4$ (Lemma 2), and recalling the convergence results of $\boldsymbol{\theta}^{f}$ from (18), we obtain the following results:
$\sum_{k=0}^{K-1}\alpha_{k}^{\mu}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}(\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}-\frac{\sum_{k=0}^{K-1}(\alpha_{k}^{\mu})^{2}L_{J\theta}^{2}L_{J\theta\theta}}{2}$ (36)
$\leq$ $\hat{J}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})+C_{1}\Big((1-\nu_{J})^{K/2}$
$\times\parallel\boldsymbol{\theta}_{0}^{J}-\boldsymbol{\theta}^{J*}\parallel+\sqrt{\frac{\nu_{f}}{2+K}}+2\sum_{k=0}^{K-1}(\sqrt{\frac{\nu_{f}}{2+k+1}}$
@@
Furthermore, applying the geometric series formula ($\sum_{k=0}^{\infty}ar^{k}=\frac{a}{1-r},|r|<1$) and the inequality $\sum_{k=1}^{K}k^{-1/2}\approx\int_{1}^{K}k^{-1/2}\leq 2\sqrt{K}$ to (36), and then taking the minimum of the left-hand side of (36), we obtain:
$\min_{k\in\{0,1,2,\cdots,K-1\}}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}(\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}$ (37)
@@
$+\sqrt{\nu_{f}}(K^{-1/2}+4K^{1/2})+\frac{1}{C_{1}}(\hat{J}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})$
$-\hat{J}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*}))\Big)+\frac{\sum_{k=0}^{K-1}(\alpha_{k}^{\mu})^{2}L_{J\theta}^{2}L_{J\theta\theta}}{2\sum_{k=0}^{K-1}\alpha_{k}^{\mu}}.$
To achieve $\lim_{k\rightarrow\infty}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}(\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}=0$, one needs to choose $\sum_{k=0}^{\infty}\alpha_{k}^{\mu}=\infty$ and $\sum_{k=0}^{\infty}(\alpha_{k}^{\mu})^{2}<\infty$.
@@
$\min_{k\in\{0,1,2,\cdots,K-1\}}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}(\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}$
@@
$+\hat{J}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})\Big)$
@@
$\quad+\hat{J}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})\Big),$